OpenStudy (anonymous):

Solve x – 2y = 4 for y.

OpenStudy (anonymous):

x-2y=4 -2y=4-x 2y=x-4 y=(x-4)/2

OpenStudy (anonymous):

x-2y=4 -x -x -2y=x-4 /-2 /-2 y=-1/2x+2 >:)